OpenStudy (anonymous):

Is x = 1 a solution of the equation 2 – 8x = –6?

sam (.sam.):

Did you solve the equation? or is that your answer

OpenStudy (anonymous):

i need help with it

sam (.sam.):

The question is asking you to prove if x=1, so, you have to solve it. 2 – 8x = –6 Subtract both sides by 2, 2-2-8x=-6-2 -8x=-8 Divide both sides by -8 x=1 So, you have proven that 'x' in 2 – 8x = -6 is equal to 1.

OpenStudy (anonymous):

thanks so much sam